How do I sove for x in this equation?
X + 12/x = 7
I know I set it equal to 0 right? So it'd be x+12/x + 7 = 0
But I'm not sure what to do with that x in the denominator. What do I do next?

Answer by scott8148(6628)

first, clear the fraction by multiplying by x

x^2 + 12 = 7x

now, set to zero ___ x^2 - 7x + 12 = 0

factoring ___ (x - 3)(x - 4) = 0

so x equals 3 and 4 (two solutions)